# Quantum Pufferfish Privacy: A Flexible Privacy Framework for Quantum Systems

### v1

@@ Abstract
Auditing privacy, privacy-utility tradeoff, pufferfish privacy, quantum differential privacy.
@@ I Introduction
Figure 1: Representation of a setup in which the goal is to hide the resource value $\mathsf{V}$ being $a$ or $b$ on the set of states $\{\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}\}$. In this abstract diagram, large squares represent the entire quantum state, while small rectangles correspond to a specific attribute of the quantum state. The specific attribute can take on one of two values, either $a$ or $b$, where $a$ is represented by solid lines and $b$ is represented by dotted lines.
With a surging interest in quantum and classical–quantum systems, ensuring privacy of both classical and quantum data has become pivotal. Privacy-preserving data analysis has been widely studied for classical systems by means of statistical privacy frameworks. Differential privacy (DP) is an important such framework, which enables answering aggregate queries about a database while keeping individual records private [1, 2]. However, DP accounts for one type of private information only (namely, records of individual users), and it does not allow encoding domain knowledge into the framework. To address these limitations, a versatile generalization of DP, termed Pufferfish Privacy (PP), has been proposed [3]. PP allows for customizing which information is regarded as private and explicitly integrates distributional assumptions into the definition [3, 4, 5]. Information-theoretic formulations of classical DP and PP have been proposed in [6] and [7], respectively.
Quantum DP (QDP) is a generalization of the classical DP notion and has been proposed in [8]. See also [9] for DP of quantum measurements and [10] for an information-theoretic interpretation of QDP. Connections to quantum stability through private learning have been studied in [11, 12]. Moreover, [13] has explored how quantum classifiers can be made private by using the intrinsic noise of existing quantum systems. However, similar to the classical case, the versatility of QDP is limited.
@@ I-A Motivation
Flexible secrets. QDP guarantees that any pair of states that are classified as neighbors are approximately indistinguishable, i.e., cannot be identified under any possible measurement. However, scenarios may arise in which one wants to hide specific properties of the states, as opposed to the state itself (e.g., whether the states possess a certain symmetry or property, have any entanglement with a special subsystem, or have secret correlations with other systems). In such situations, QDP may be an overly pessimistic notion of privacy, which, in turn, hinders utility. As an example, consider hiding a resource value $\mathsf{V}$ of the states $\{\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}\}$, for which $\mathsf{V}(\rho_{1})=\mathsf{V}(\rho_{2})=a$ and $\mathsf{V}(\sigma_{1})=\mathsf{V}(\sigma_{2})=b$. As illustrated in Fig. 1, hiding whether $\mathsf{V}$ equals $a$ or $b$ amounts to making the classes $\{\rho_{1},\rho_{2}\}$ and $\{\sigma_{1},\sigma_{2}\}$ indistinguishable. This can be achieved by applying a QDP mechanism to the state space $\{\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}\}$, but doing so provides a stricter guarantee than required. The source of the issue is the inability of QDP to account for secrets concerning collections of states (as opposed to singletons), which is the first issue we aim to address.
@@ I-B Contributions
We then move on to derive properties of QPP mechanisms, encompassing convexity, post-processing, and composability (both parallel and adaptive). As a specific example, we characterize the flip parameter that guarantees QPP of the depolarization mechanism. We also describe how QPP mechanisms implementable on quantum devices can be instantiated to achieve classical PP. We consider the associated privacy-utility tradeoff for QPP mechanisms. Our utility metric captures how invertible the privacy mechanism is, which is formulated as the infimized diamond distance between a post-processing of the mechanism’s output and the identity channel. We show that this utility metric can be computed as an SDP and analyze the privacy-utility tradeoff of the depolarization mechanism. Lastly, we study optimal privacy-utility tradeoffs of QPP mechanisms and characterize the achievable region in several settings.
@@ II-A Notation
Sets are denoted by calligraphic letters, e.g., $\mathcal{X}$. For $k,n\in\mathbb{N}$, we use $\mathcal{X}^{n\times k}$ to denote the database space of $n\times k$ matrices; columns correspond to different attributes while rows to different individuals. The $(i,j)$th entry of $x\in\mathcal{X}^{n\times k}$ is denoted as $x(i,j)$. The $i$th row and $j$th column of $x$ are denoted by $x(i,\cdot)$ and $x(\cdot,j)$, respectively.
@@
We denote by $(\Omega,\mathcal{F},\mathbb{P})$ the underlying probability space on which all random variables (RVs) are defined, with $\mathbb{E}$ designating expectation. RVs are denoted by upper case letters, e.g., $X$, with $P_{X}$ representing the corresponding probability law. For $X\sim P_{X}$, we interchangeably use $\mathrm{supp}(X)$ and $\mathrm{supp}(P_{X})$ for the support. The joint law of $(X,Y)$ is denoted by $P_{XY}$, while $P_{Y|X}$ designates the (regular) conditional probability of $Y$ given $X$. Conventions for $n\times k$-dimensional random variables are the same as for deterministic elements. The space of all Borel probability measures on $\mathcal{S}\subseteq\mathbb{R}^{d}$ is denoted by $\mathcal{P}(\mathcal{S})$. The Kullback–Leibler (KL) divergence between $P,Q\in\mathcal{P}(\mathcal{X})$ with $P\ll Q$ is given by $\mathsf{D}(P\|Q)\coloneqq\mathbb{E}_{P}\!\left[\ln\!\left(\frac{dP}{dQ}\right)\right],$ where $\frac{dP}{dQ}$ is the Radon–Nikodym derivative of $P$ with respect to $Q$. For $(X,Y)\sim P_{XY}$, the mutual information between $X$ and $Y$ is denoted by $\mathsf{I}(X;Y)\coloneqq\mathsf{D}(P_{XY}\|P_{X}\otimes P_{Y})$.
@@
We now review basic concepts from quantum information theory and refer the reader to [26, 27] for more details. A (classical or quantum) system $R$ is identified with a finite-dimensional Hilbert space $\mathcal{H}_{R}$. We denote the set of linear operators acting on $\mathcal{H}_{R}$ by $\mathcal{L}(\mathcal{H}_{R})$. The support of a linear operator $X\in\mathcal{L}(\mathcal{H}_{R})$ is defined to be the orthogonal complement of its kernel, and it is denoted by $\mathrm{supp}(X)$. Let $\mathrm{T}(C)$ denote the transpose of $C$. The partial transpose of $C\in\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ on the subsystem $A$ is represented as $\mathrm{T}_{A}(C)$. Let $\mathrm{Tr}\!\left[C\right]$ denote the trace of $C$, and let $\mathrm{Tr}_{A}\!\left[C\right]$ denote the partial trace of $C$ over the subsystem $A$. The trace norm of a matrix $B$ is defined as $\left\|B\right\|_{1}\coloneqq\mathrm{Tr}[\sqrt{B^{\dagger}B}]$. For operators $A$ and $B$, the notation $A\geq B$ indicates that $A-B$ is a PSD operator, while $A>B$ indicates that $A-B$ is a positive definite operator.
@@
A quantum state $\rho_{R}\in\mathcal{L}(\mathcal{H}_{R})$ on $R$ is a positive semi-definite (PSD), unit-trace operator acting on $\mathcal{H}_{R}$. We denote the set of all density operators in $\mathcal{L}(\mathcal{H}_{R})$ by $\mathcal{D}(\mathcal{H}_{R})$. A state $\rho_{R}$ of rank one is called pure, and we may choose a normalized vector $|\psi\rangle\in\mathcal{H}_{R}$ satisfying $\rho_{R}=|\psi\rangle\!\langle\psi|$ in this case. Otherwise, $\rho_{R}$ is called a mixed state. By the spectral decomposition theorem, every mixed state can be written as a convex combination of pure, orthogonal states. A quantum channel $\mathcal{N}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ is a linear, completely-positive and trace-preserving (CPTP) map from $\mathcal{L}(\mathcal{H}_{A})$ to $\mathcal{L}(\mathcal{H}_{B})$. We denote the adjoint of $\mathcal{N}$ by $\mathcal{N}^{\dagger}$. A measurement of a quantum system $R$ is described by a positive operator-valued measure (POVM) $\{\mathrm{M}_{y}\}_{y\in\mathcal{Y}}$, which is defined to be a collection of PSD operators satisfying $\sum_{y\in\mathcal{Y}}\mathrm{M}_{y}=\mathrm{I}_{\mathcal{H}_{R}}$, where $\mathcal{Y}$ is a finite alphabet. The Born rule dictates that, after applying the above POVM to $\rho\in\mathcal{D}(\mathcal{H}_{R})$, the probability of observing the outcome $y$ is given by $\mathrm{Tr}\!\left[\mathrm{M}_{y}\rho\right]$.
@@ II-B Quantum Divergences
We define several quantum divergences that will be used throughout this work. We call a distinguishability measure $\mathsf{D}(\cdot\|\cdot)$ a generalized divergence [28] if it satisfies the data-processing inequality; i.e., for every channel $\mathcal{N}$, state $\rho$, and PSD operator $\sigma$,
@@
where the optimization in the definition is over every bipartite density operator $\rho_{RA}$, with an arbitrarily large system $R$. It is well known, however, that it suffices to perform the optimization over pure bipartite states such that the dimension of the reference system $R$ is equal to the dimension of the channel input system $A$.
@@
$\mathsf{D}_{\alpha}(\rho\|\sigma)\coloneqq\\ \begin{cases}\frac{1}{\alpha-1}\ln\mathrm{Tr}[\rho^{\alpha}\sigma^{1-\alpha}],&\mbox{if }\alpha\in(0,1)\lor\\ &(\alpha>1\land\mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma))\\ +\infty,&\mbox{otherwise}\end{cases}$ (5)
The special case of $\alpha\to 1$ is called the quantum relative entropy and amounts to
@@
Equivalently, $\mathsf{S}(\rho)=-\mathsf{D}_{1}(\rho\|I)$, where $I$ is the identity operator.
@@
$\widetilde{D}_{\alpha}(\rho\|\sigma)\coloneqq\\ \begin{cases}\frac{1}{\alpha-1}\ln\widetilde{Q}_{\alpha}(\rho\|\sigma)&\mbox{if}\ \alpha\in(0,1)\,\lor\\ &(\alpha\in(1,\infty)\land\ \mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma)),\\ +\infty&\mbox{otherwise},\end{cases}$ (8)
where
$\widetilde{Q}_{\alpha}(\rho\|\sigma)\coloneqq\mathrm{Tr}\!\left[\left(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}}\right)^{\alpha}\right].$ (9)
It is a generalized divergence for $\alpha\in[1/2,1)\cup(1,\infty)$ [35] (see also [36]).
@@
and so we can speak of a single DL divergence, which we set hereafter to be $\overline{\mathsf{D}}^{\delta}$ from (10b). Slightly rewriting (II-B), we have the equivalent representations:
@@
The latter forms are helpful for making connections to privacy.
@@
$=\log\sup_{0\leq\mathrm{M}\leq\mathrm{I}}\frac{\mathrm{Tr}\!\left[\mathrm{M}\rho\right]}{\mathrm{Tr}\!\left[\mathrm{M}\sigma\right]},$ (15)
@@ II-C Classical and Quantum Privacy
In this section, we provide background on the existing definitions of privacy for both classical and quantum systems, starting from classical DP and proceeding to quantum DP.
@@ II-C1 Classical Differential and Pufferfish Privacy
DP allows for answering queries about aggregate quantities while protecting the individual entries in a database [1]. To this end, the output of a differentially privacy mechanism should be indistinguishable for neighboring databases, defined as those that differ only in a single record (row). Formally, we say that $x,x^{\prime}\in\mathcal{X}^{n\times k}$ are neighbors, denoted $x\sim x^{\prime}$, if $x(i,\cdot)\neq x^{\prime}(i,\cdot)$ for some $i\in\{1,\ldots,n\}$, and they agree on all other rows. We also note that a randomized privacy mechanism $A$, as mentioned below, is described by a (regular) conditional probability distribution $P_{A|X}$ for its output given the data.
@@
for all $P_{X}\in\Theta$, $(\mathcal{R},\mathcal{T})\in\mathcal{Q}_{c}$ with $P_{X}(\mathcal{R}),P_{X}(\mathcal{T})>0$, and $\mathcal{B}\subseteq\mathcal{Y}$ measurable.

### v2

@@ Abstract
Auditing privacy, privacy-utility tradeoff, pufferfish privacy, quantum differential privacy, quantum generalized divergences
@@ I Introduction
With a surging interest in quantum and hybrid classical–quantum systems, ensuring privacy of both classical and quantum data has become pivotal. Privacy-preserving data analysis has been widely studied for classical systems by means of statistical privacy frameworks. Differential privacy (DP) is an important statistical privacy framework that enables answering aggregate queries about a database while keeping individual records private [1, 2]. However, DP accounts for one type of private information only (namely, records of individual users), and it does not allow encoding domain knowledge into the framework. To address these limitations, a versatile generalization of DP, termed Pufferfish Privacy (PP), has been proposed [3]. PP allows for customizing which information is regarded as private and explicitly integrates distributional assumptions into the definition [3, 4, 5]. PP has found use in several applications, including smart metering [6, 7] and trajectory monitoring with location tracking [8, 9] (see also Figure 1 of [10] for an explicit example related to salary releases, where PP is applicable). Information-theoretic formulations of classical DP and PP have been proposed in [11] and [12], respectively.
Quantum DP (QDP) is a generalization of the classical DP notion and has been proposed in [13]. See also [14] for DP of quantum measurements and [15] for an information-theoretic interpretation of QDP. Connections to quantum stability through private learning have been studied in [16]. Moreover, [17] has explored how quantum classifiers can be made private by using the intrinsic noise of existing quantum systems. See also [18, 19, 20, 21, 22] for applications of DP in quantum machine learning. Additionally, privacy amplification of quantum and quantum inspired algorithms has been analysed using QDP and classical DP notions in [23]. However, similar to the classical case, the versatility of QDP is limited.
@@ I-A Motivation
Figure 1: Depiction of a setup where the goal is to hide whether the amount entanglement $\mathsf{V}$ present in the bipartite states $\rho_{1}$, $\rho_{2}$, $\sigma_{1}$, and $\sigma_{2}$ equals $a$ or $b$. In this diagram, large squares represent the entire quantum state, while small rectangles correspond to a specific attribute of that state (i.e., the amount of entanglement as quantified by the function $\mathsf{V}$). The specific attribute can take on one of two values, $a$ or $b$, represented by solid or dotted lines, respectively. As the goal is to conceal only the entanglement level, and not necessarily the specific quantum state, we want the sets =$\mathcal{R}=\{\rho_{1},\rho_{2}\}$ and $\mathcal{T}=\{\sigma_{1},\sigma_{2}\}$ to be indistinguishable.
@@
Flexible secrets. QDP guarantees that any pair of states that are classified as neighbors are approximately indistinguishable, i.e., cannot be identified under any possible measurement. However, scenarios may arise in which one wants to hide specific properties of the states, as opposed to the state itself (e.g., whether the states possess a certain symmetry or property, have any entanglement with a special subsystem, or have secret correlations with other systems). In such situations, QDP may be an overly pessimistic notion of privacy, which, in turn, hinders utility. As an example, consider hiding the amount of entanglement $\mathsf{V}$ present in the bipartite states in the set $\{\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}\}$, for which $\mathsf{V}(\rho_{1})=\mathsf{V}(\rho_{2})=a$ and $\mathsf{V}(\sigma_{1})=\mathsf{V}(\sigma_{2})=b$. As illustrated in Fig. 1, hiding whether $\mathsf{V}$ equals $a$ or $b$ amounts to making the classes $\{\rho_{1},\rho_{2}\}$ and $\{\sigma_{1},\sigma_{2}\}$ indistinguishable. This can be achieved by applying a QDP mechanism to the state space $\{\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}\}$ by choosing $(\rho_{i},\sigma_{j})$ for all $i,j\in\{1,2\}$ as neighbors, with the criterion that $\rho$ and $\sigma$ are neighbors if and only if $|\mathsf{V}(\rho)-\mathsf{V}(\sigma)|=|a-b|$. However, doing so provides a stricter guarantee than required. The source of the issue is the inability of QDP to account for secrets concerning collections of states (as opposed to singletons), which is the first issue we aim to address.
@@ I-B Contributions
We then move on to derive properties of QPP mechanisms, encompassing convexity, post-processing, and composability (both parallel and adaptive). As a specific example, we characterize the flip parameter that guarantees QPP of the depolarization mechanism. We also describe how QPP mechanisms implementable on quantum devices can be instantiated to achieve classical PP. We consider the associated privacy-utility tradeoff for QPP mechanisms. Our utility metric captures how invertible the privacy mechanism is, which is formulated as the infimized diamond distance between a post-processing of the mechanism’s output and the identity channel. We show that this utility metric can be computed as an SDP, and we analyze the privacy-utility tradeoff of the depolarization mechanism. Lastly, we study optimal privacy-utility tradeoffs of QPP mechanisms and characterize the achievable region in several settings.
@@ II-A Notation
Sets are denoted by calligraphic letters, e.g., $\mathcal{X}$. For $k,n\in\mathbb{N}$, we use $\mathcal{X}^{n\times k}$ to denote the database space of $n\times k$ matrices; columns correspond to different attributes while rows to different individuals. The $(i,j)$th entry of $x\in\mathcal{X}^{n\times k}$ is denoted as $x(i,j)$. The $i$th row and $j$th column of $x$ are denoted by $x(i,\cdot)$ and $x(\cdot,j)$, respectively. We denote by $(\Omega,\mathcal{F},\mathbb{P})$ the underlying probability space on which all random variables (RVs) are defined, with $\mathbb{E}$ designating expectation. RVs are denoted by upper case letters, e.g., $X$, with $P_{X}$ representing the corresponding probability law. For $X\sim P_{X}$, we interchangeably use $\mathrm{supp}(X)$ and $\mathrm{supp}(P_{X})$ for the support. The joint law of $(X,Y)$ is denoted by $P_{XY}$, while $P_{Y|X}$ designates the (regular) conditional probability of $Y$ given $X$. Conventions for $n\times k$-dimensional random variables are the same as for deterministic elements. The space of all Borel probability measures on $\mathcal{S}\subseteq\mathbb{R}^{d}$ is denoted by $\mathcal{P}(\mathcal{S})$. The Kullback–Leibler (KL) divergence between $P,Q\in\mathcal{P}(\mathcal{X})$ with $P\ll Q$ is given by $\mathsf{D}(P\|Q)\coloneqq\mathbb{E}_{P}\!\left[\ln\!\left(\frac{dP}{dQ}\right)\right],$ where $\frac{dP}{dQ}$ is the Radon–Nikodym derivative of $P$ with respect to $Q$. For $(X,Y)\sim P_{XY}$, the mutual information between $X$ and $Y$ is denoted by $\mathsf{I}(X;Y)\coloneqq\mathsf{D}(P_{XY}\|P_{X}\otimes P_{Y})$.
@@
We now review basic concepts from quantum information theory and refer the reader to [36, 37] for more details. A (classical or quantum) system $R$ is identified with a finite-dimensional Hilbert space $\mathcal{H}_{R}$. We denote the set of linear operators acting on $\mathcal{H}_{R}$ by $\mathcal{L}(\mathcal{H}_{R})$. The support of a linear operator $X\in\mathcal{L}(\mathcal{H}_{R})$ is defined to be the orthogonal complement of its kernel, and it is denoted by $\mathrm{supp}(X)$. Let $\mathrm{T}(C)$ denote the transpose of $C$. The partial transpose of $C\in\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ on the subsystem $A$ is represented as $\mathrm{T}_{A}(C)$. Let $\mathrm{Tr}\!\left[C\right]$ denote the trace of $C$, and let $\mathrm{Tr}_{A}\!\left[C\right]$ denote the partial trace of $C$ over the subsystem $A$. The trace norm of a matrix $B$ is defined as $\left\|B\right\|_{1}\coloneqq\mathrm{Tr}[\sqrt{B^{\dagger}B}]$. For operators $A$ and $B$, the notation $A\geq B$ indicates that $A-B$ is a positive semi-definite (PSD) operator, while $A>B$ indicates that $A-B$ is a positive definite operator.
@@
A quantum state $\rho_{R}\in\mathcal{L}(\mathcal{H}_{R})$ on $R$ is a PSD, unit-trace operator acting on $\mathcal{H}_{R}$. We denote the set of all density operators in $\mathcal{L}(\mathcal{H}_{R})$ by $\mathcal{D}(\mathcal{H}_{R})$. A state $\rho_{R}$ of rank one is called pure, and we may choose a normalized vector $|\psi\rangle\in\mathcal{H}_{R}$ satisfying $\rho_{R}=|\psi\rangle\!\langle\psi|$ in this case. Otherwise, $\rho_{R}$ is called a mixed state. By the spectral decomposition theorem, every mixed state can be written as a convex combination of pure, orthogonal states. A quantum channel $\mathcal{N}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ is a linear, completely positive and trace-preserving (CPTP) map from $\mathcal{L}(\mathcal{H}_{A})$ to $\mathcal{L}(\mathcal{H}_{B})$. We denote the adjoint of $\mathcal{N}$ by $\mathcal{N}^{\dagger}$. A measurement of a quantum system $R$ is described by a positive operator-valued measure (POVM) $\{\mathrm{M}_{y}\}_{y\in\mathcal{Y}}$, which is defined to be a collection of PSD operators satisfying $\sum_{y\in\mathcal{Y}}\mathrm{M}_{y}=\mathrm{I}_{\mathcal{H}_{R}}$, where $\mathcal{Y}$ is a finite alphabet. The Born rule dictates that, after applying the above POVM to $\rho\in\mathcal{D}(\mathcal{H}_{R})$, the probability of observing the outcome $y$ is given by $\mathrm{Tr}\!\left[\mathrm{M}_{y}\rho\right]$.
@@ II-B Quantum Divergences
We define several quantum divergences that will be used throughout this work. We call a distinguishability measure $\boldsymbol{\mathsf{D}}(\cdot\|\cdot)$ a generalized divergence [38] if it satisfies the data-processing inequality; i.e., for every channel $\mathcal{N}$, state $\rho$, and PSD operator $\sigma$,
@@
where the optimization in the definition is over every reference system $R$ and bipartite density operator $\rho_{RA}$ (with $R$ allowed to be arbitrarily large). It is well known, however, that it suffices to perform the optimization over pure bipartite states such that the dimension of the reference system $R$ is equal to the dimension of the channel input system $A$.
@@
$\mathsf{D}_{\alpha}(\rho\|\sigma)\coloneqq\frac{1}{\alpha-1}\ln\mathrm{Tr}[\rho^{\alpha}\sigma^{1-\alpha}]$ (5)
if $\alpha\in(0,1)\lor(\alpha>1\land\mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma))$ and $\infty$ otherwise. It is a generalized divergence for $\alpha\in[0,1)\cup(1,2]$ [42]. The special case of $\alpha\to 1$ is called the quantum relative entropy and amounts to
@@
Equivalently, $\mathsf{S}(\rho)=-\mathsf{D}_{1}(\rho\|\mathrm{I})$, where $\mathrm{I}$ is the identity operator.
@@
$\widetilde{\mathsf{D}}_{\alpha}(\rho\|\sigma)\coloneqq\frac{1}{\alpha-1}\ln\mathrm{Tr}\!\left[\left(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}}\right)^{\alpha}\right]$ (8)
if $\alpha\in(0,1)\,\lor(\alpha\in(1,\infty)\land\ \mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma))$ and $\infty$ otherwise. It is a generalized divergence for $\alpha\in[1/2,1)\cup(1,\infty)$ [45] (see also [46]).
@@
and so we can speak of a single DL divergence, which we set hereafter to be $\overline{\mathsf{D}}^{\delta}$ from (9b). Slightly rewriting (9), we have the equivalent representations:
@@
$=\ln\sup_{0\leq\mathrm{M}\leq\mathrm{I}}\frac{\mathrm{Tr}\!\left[\mathrm{M}\rho\right]}{\mathrm{Tr}\!\left[\mathrm{M}\sigma\right]},$ (14)
@@ II-C Classical and Quantum Privacy
In this section, we provide background on the existing definitions of privacy for both classical and quantum systems, starting from classical DP and proceeding to quantum DP thereafter.
@@ II-C1 Classical Differential and Pufferfish Privacy
DP allows for answering queries about aggregate quantities while protecting the individual entries in a database [1]. To this end, the output of a differential privacy mechanism should be indistinguishable for neighboring databases, defined as those that differ only in a single record (row). Formally, we say that $x,x^{\prime}\in\mathcal{X}^{n\times k}$ are neighbors, denoted $x\sim x^{\prime}$, if $x(i,\cdot)\neq x^{\prime}(i,\cdot)$ for some $i\in\{1,\ldots,n\}$, and they agree on all other rows. We also note that a randomized privacy mechanism $A$, as mentioned below, is described by a (regular) conditional probability distribution $P_{A|X}$ for its output given the data.
@@
for all $P_{X}\in\Theta$, $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ with $P_{X}(\mathcal{R}),P_{X}(\mathcal{T})>0$, and $\mathcal{B}\subseteq\mathcal{Y}$ measurable.